OpenStudy (anonymous):

MEDAL TO BEST ANSWER!!! x2 - 8x + 12 Factor this quadratic and show all work 4x2 - 6x + 2 = 0 Solve for x and show all work

OpenStudy (anonymous):

Please help!!!!!

OpenStudy (anonymous):

so you always have to think like this: what are the number added and give the middle (-8x) and also the product gives 12. can you think of any?

OpenStudy (anonymous):

for the first one be (x-4)(x-3)? @Supenova

OpenStudy (anonymous):

that will give you x2-7x+12

OpenStudy (anonymous):

does -4 X (-3)= 12?

OpenStudy (anonymous):

and now u can try to do the next one

OpenStudy (anonymous):

yes. a negative times a negative equals a positive @hello1213

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but what is ur answer?

OpenStudy (anonymous):

it has to be like this 4x2-4x-2x+2 . u see now u can factor

OpenStudy (anonymous):

by distributive property

OpenStudy (anonymous):

4x(x-1)-2(x-1) so u have (4x-2)(x-1)

OpenStudy (anonymous):

x2 -8x +12 break this down to x2 -6x -2x +12 ( think of two negative numbers which multiply to give 12 and add to give -8, which in this case are -6 and -2) taking like terms out : x(x-6) -2(x-6) you see that the terms in brackets are the same therefore : (x-2)(x-6)

OpenStudy (anonymous):

4x-2 and x-1

OpenStudy (anonymous):

u are close

OpenStudy (anonymous):

and on the first one Ralph's answer is right

OpenStudy (anonymous):

u r right too @smileandcanter

OpenStudy (anonymous):

when u factor it you should have x-2 and x-6. mulitplying these two should give u x2-8x+12

OpenStudy (anonymous):

4x(x-1)-2(x-1) take the like terms and common take only one common factor of (x-1) (4x-2)(x-1)

OpenStudy (anonymous):

for the second one : 4x2 -6x +2 =0 multiply the first and last constants together ( i.e multiply 4 and 2) The product is 8. Think of two numbers which multiply to give you 8 and add up to give you the middle constant (-6). That would be -4 and -2 so rewrite the equation : 4x2 -4x -2x +2 =0 factorize: 4x(x-1) -2(x-1) =0 (4x-2)(x-1)=0 one of the terms has to be zero if the product of the terms is zero. therefore 4x-2=0 or x-1=0 so, x=0.5 or x=1